OpenStudy (anonymous):

Find w(6) for the function w(x) = 4x + 5. 29 15 44 51

OpenStudy (anonymous):

29

OpenStudy (anonymous):

4 multiplied by 6 then take that total and multiple it by 5

jimthompson5910 (jim_thompson5910):

w(x) = 4x + 5 w(6) = 4(6) + 5 ... replace each x with 6, then evaluate (using PEMDAS) w(6) = 24 + 5 w(6) = 29